OpenStudy (toxicsugar22):

Ronald corp uses a perpetual inventory system. determine the costs assigned to cost of goods sold and ending inventory using both the FIFO and LIFO methods.

OpenStudy (toxicsugar22):

date Activities units acquired at cost unit sold 1/1 Beginning inventory 100 units @ $10= $1000 1/10 Sales 90 units 3/14 Purchase 250 units @ $15= 3750 3/15 Sales 140 units 7/30 Purchase 400 units @ $20= 8000 10/5 Sales 300 units 10/26 Purchase 600 units @ $25= 15000 Totals 1350 units $27750 530 units

OpenStudy (theeric):

I imagine LIFO = last in first out and FIFO = first in first out, but I don't know about this.

OpenStudy (theeric):

Good luck!

OpenStudy (anonymous):

FIFO stands for first-in, first-out, meaning that the oldest inventory items are recorded as sold first LIFO stands for last-in, first-out, meaning that the most recently produced items are recorded as sold first

OpenStudy (anonymous):

Errm I'm not really sure how to do this , sorry :(

OpenStudy (anonymous):

@Whitemonsterbunny17 @halorazer @Drsuz98 @Mr_Perfection_xD can u help this young lady out

whitemonsterbunny17 (whitemonsterbunny17):

I don't know how to solve this type of problem, sorry. :/ I wish I could've been more helpful!! :s

OpenStudy (mr_perfection_xd):

I'm so sorry. :I but all of this is greek to me. what class is this? im sophomore btw.

OpenStudy (toxicsugar22):

This is accounting

OpenStudy (anonymous):

i believe in this instance FIFO and LIFO are referring to the sequence in which the orders are filled, and I think that the wrapping of the displayed table is making it look more confusing than it actually is.